### v1

# Understanding Self-Regulated Learning Behavior Among High and Low Dropout Risk Students During CS1: Combining Trace Logs, Dropout Prediction and Self-Reports

Abstract.
The introductory programming course (CS1) at the university level is often perceived as particularly challenging, contributing to high dropout rates among Computer Science students. Identifying when and how students encounter difficulties in this course is critical for providing targeted support. This study explores the behavioral patterns of CS1 students at varying dropout risks using self-regulated learning (SRL) as the theoretical framework. Using learning analytics, we analyzed trace logs and task performance data from a virtual learning environment to map resource usage patterns and used student dropout prediction to distinguish between low and high dropout risk behaviors. Data from 47 consenting students were used to carry out the analysis. Additionally, self-report questionnaires from 29 participants enriched the interpretation of observed patterns. The findings reveal distinct weekly learning strategy types and categorize course behavior. Among low dropout risk students, three learning strategies were identified that different in how students prioritized completing tasks and reading course materials. High dropout risk students exhibited nine different strategies, some representing temporary unsuccessful strategies that can be recovered from, while others indicating behaviors of students on the verge of dropping out. This study highlights the value of combining student behavior profiling with predictive learning analytics to explain dropout predictions and devise targeted interventions. Practical findings of the study can in turn be used to help teachers, teaching assistants and other practitioners to better recognize and address students at the verge of dropping out.
self-regulation, CS1, trace logs, student dropout prediction
copyright: ccccs: Social and professional topics CS1
1. Introduction
Developing solid learning and working skills is paramount when transitioning to higher education. When students begin their university studies, they may face numerous challenges, such as increased workload, higher expectations, and greater emphasis on independent learning (Leese, 2010). This notion is especially true in Computer Science, where the introductory programming course (CS1) is often perceived as laborious and fast-paced by first-year students (e.g., Hämäläinen and Isomöttönen, 2019). Students who struggle with the learning process may face hits to their self-efficacy beliefs, affecting later studies and potentially leading to burnout (e.g., Järvinen et al., 2024). Lack of motivation, perceived course difficulty, and low interest can, at worst, result in dropping out (Petersen et al., 2016; Barr and Kallia, 2022). To mitigate dropout rates and foster effective learning habits, universities employ various support mechanisms, such as introductory courses, peer mentoring programs and course-specific interventions (e.g., Turner et al., 2017; Kangas et al., 2017; Pon-Barry et al., 2017; Borrella et al., 2022).
Identifying struggling students and those at risk of dropping out is crucial to ensuring these interventions are effective. Academic performance prediction and student dropout detection have been extensively researched (Hellas et al., 2018; Prenkaj et al., 2020). Many studies apply learning analytics techniques to various student data, including demographics (e.g., Naseem et al., 2019), basic activity indicators (e.g., Waheed et al., 2023), outcomes of student assignments (e.g., Zhidkikh et al., 2024) and self-reports (e.g., Tempelaar et al., 2018). Notably, Moreno-Marcos et al. (2020) evaluated an effective prediction approach using students’ self-regulation behavior data collected from self-reports and trace logs. Although most student dropout prediction methods demonstrate high accuracy, easy explainability of the predictions remains a challenge (Zhidkikh et al., 2024).
To increase the effectiveness of the predictions and interventions, it is therefore essential to understand how students study and how they utilize available learning resources. Recent research has increasingly focused on identifying student learning behaviors, emphasizing the self-regulation, metacognition, and motivation aspects (Lorås et al., 2022). For instance, common activity indicators can be used to model students’ learning approaches and to identify student behavior profiles (e.g., Kim et al., 2020; Sharma et al., 2015). Similarly, trace logs from virtual learning environments can reveal patterns in students’ learning resource usage and detect common learning strategies used in a classroom (e.g., Jovanović et al., 2017; Zhidkikh et al., 2023). Additionally, self-reports, such as questionnaires, learning journals, and interviews may provide insights into students’ use of learning strategies and cognitive processes (e.g., Loksa and Ko, 2016; Loksa et al., 2020; Silva et al., 2024). However, while these methods help map learning behaviors, they usually do not directly link behavior to early dropout (Zhidkikh et al., 2023). At most, some studies have compared behaviors of high and low-performing students, often in a post-course setting. For instance, Liao et al. (2019) interviewed 19 CS1 students who took the final CS1 exam and identified general differences in learning paths between higher-performing and lower-performing students.
Motivated by the examples above, this paper aims to deepen the understanding of learning behaviors among high and low-risk students during CS1. The present paper complements prior key studies (Liao et al., 2019; Moreno-Marcos et al., 2020). Specifically, we map students’ learning paths, extending previous work (cf. Liao et al., 2019) by tracking weekly learning behaviors using trace logs collected during the course. Further, we employ dropout prediction to identify at-risk students for every course week (cf. Moreno-Marcos et al., 2020) and use this information to discern weekly learning behavior by the associated dropout risk. To this end, the paper presents a learning analytics-based method to link learning resource usage patterns with dropout predictions, combining tools developed in prior studies (Zhidkikh et al., 2023, 2024). A post-course self-report questionnaire was also administered to support the interpretation of the behavioral patterns and to evaluate the potential of the learning analytics used. Both qualitative and quantitative data were inductively explored to gain insights into students’ behaviors at different dropout risk levels. We use self-regulated learning as the foundational theoretical framework widely applied in CSEd research to understand the learning process (Prather et al., 2020). The study ultimately seeks to answer the following research questions:
  1. RQ1
    What learning behaviors emerge in students when identified as low dropout risk?
  2. RQ2
    What learning behaviors emerge in students when identified as high dropout risk?
  3. RQ3
    How did students’ self-reported views on their learning inform the interpretation of behaviors from trace log and dropout prediction data?
2. Theoretical framework and related work
In this paper, we employ self-regulated learning (SRL) theory as the primary framework for mapping and understanding student behavior in CS1. As such, we first provide a brief overview into SRL and how it is measured in academic contexts. To contextualize the research and aid interpretation of the results, we also present recent related work on measuring SRL in CSEd. Finally, we present related works on student dropout prediction to provide further context on the methods applied in this study.
2.1. Self-regulated learning and learning strategies
Self-regulation generally refers to the ability to monitor and manage one’s effort, time, motivation and behavior when learning or problem-solving (Pintrich, 2000; Schunk, 2005). While metacognition describes student’s knowledge to self-evaluate learning outcomes and approaches, self-regulation describes the ability to control them (Prather et al., 2020). Therefore, SRL is a strategic learning process in which a student self-regulates by actively setting goals, engaging with the learning resources using various strategies, monitoring progress and refining their approach based on the outcomes and external learning context (Zimmerman, 2008; Winne and Hadwin, 1998; Winne and Marzouk, 2019). Although multiple models detailing the SRL process exist (Panadero, 2017; Loksa et al., 2022), SRL is often described as a cyclical process consisting of forethought and goal-setting, enacting on the goals and self-controlling, and self-reflection on the outcomes and the process (Zimmerman and Moylan, 2009). While the general SRL process can be modelled, the specific ways a student regulates their behavior can depend highly on the context, such as the subject, student’s prior experiences, the task at hand and even other students (Ben-Eliyahu and Bernacki, 2015).
Various measures of SRL have been applied in the academic contexts. Common measures include self-report tools that assess student’s aptitudes for self-regulation (e.g., questionnaires, think-aloud protocols, learning journals), and event measures that map SRL processes via observable, externally collected data (Roth et al., 2016). While some researchers have noted issues related to the reliability of self-reports compared to event-based measures (e.g., Choi et al., 2023; Zimmerman, 2008), a growing number of studies has argued for the combined use of both measures (e.g., Fan et al., 2023; Lim et al., 2021; Tempelaar et al., 2020). As Roth et al. (2016) summarize, the triangulation of new and established measures enables researchers to gain new insights into SRL.
A common aspect used to evaluate and study SRL is the regulation of behavior through learning strategies. A learning strategy is generally a set of planned actions, such as rehearsing, seeking assistance and reviewing course materials, aimed to acquire information and skills relevant to the goal or task at hand (Zimmerman and Martinez-Pons, 1986). For event-based measures, a more fine-grained description of learning strategies, as proposed by Winne (2001) and elaborated by Winne and Marzouk (2019), can be used. In their model, student behavior can be described through simple, atomic actions, such as reading a text section or writing a note. A sequence of actions aimed at completing a specific sub-goal or a task comprises a learning tactic. For example, a student may first aim to get basic understanding of a new programming topic by perusing parts of the coursebook, reading and completing examples within the book — all as part of a single coursebook-oriented tactic. To reach the learning goals, students may adjust their tactics over time based on newly set sub-goals, self-evaluations and changing learning context. Therefore, a student’s learning strategy comprises a set of learning tactics and their approach to varying them to reach a larger learning goal. For instance, to complete programming exercises for a week, a student’s strategy may be to first use a coursebook-oriented tactic to get more content knowledge, then upon successful understanding shift to more task-oriented tactic which encompasses completing tasks while using course materials as a reference. Overall, such event-based characterization of learning strategies enables the description of the study process as an enactment of learning actions based on the current conditions. Consequently, this definition allows for the graphical representation of student behavior as sequences of actions, tactics, and strategies over time (e.g., Hadwin et al., 2007).
2.2. Self-regulated learning in computing education
There have been several prior studies on self-regulation in computing education. Collectively, many studies highlight the importance of SRL in computing education, and especially the act of supporting SRL for novice learners. The findings across many of such studies underscore the role of SRL in academic success, particularly in computer science and programming education. For example, Schraw et al. (2006) identified cognition, metacognition, and motivation as the core components of SRL and recommended strategies to promote self-regulation in science education. Falkner et al. (2014) identified discipline-specific SRL strategies that are especially useful in software development courses, emphasizing the need for cognitive and metacognitive scaffolding. Cigdem (2015) also shows that self-regulation positively influences programming achievement, particularly when coupled with high self-efficacy in blended learning environments.
In programming education, studies like Loksa and Ko (2016) and Silva et al. (2024) have shown that self-regulation strategies such as planning, monitoring, and reflection are linked to problem-solving success, although their effectiveness depends on programming skill. Along similar lines, and with those of Schraw et al. (2006), Bergin et al. (2005) found that high-performing students use more metacognitive and resource management strategies alongside intrinsic motivation to enhance their performance, supporting the need for targeted SRL training as advocated by Bielaczyc et al. (1995) and Silva et al. (2024). In contrast, Flanigan et al. (2023) found that poor SRL experiences in undergraduate computer science students can reinforce a fixed mindset. That is, fostering effective SRL strategies is essential for encouraging a growth mindset, but a failed attempt to support SRL is not detrimental only on applying SRL, but also on course performance. Finally, Liao et al. (2019) reported on multiple strategies applied by high-performing programming students such as encouraging consistent study habits, active help-seeking, reflective practices, timely feedback, and creating a supportive learning environment.
Silva et al. (2024) also discussed the challenges students face in maintaining motivation and managing time, which reinforces the need for educational interventions to help students better apply SRL strategies. Jovanović et al. (2017) found that students in flipped classrooms often rely on less effective, passive strategies, such as video watching, rather than more active approaches like formative assessments. Similarly, Matcha et al. (2019) and Liao et al. (2019) discussed how different analytic approaches can identify varying learning strategies in online courses, and argued for selecting appropriate methods such as tailoring and scaffolding to better support learners. Similar insights were also discussed by Panadero (2017) who reviewed six SRL models, concluding that these frameworks should be tailored to students’ developmental stages. Focusing their review of SRL models applied in computing education, Loksa et al. (2022) further suggested that researchers should select a framework and concepts that best fit their research context. Greene and Azevedo (2007) discussed the complexity of SRL, advocating for research that considers the interactions between cognitive and task conditions across learning phases.
Ko and Leu (2016) demonstrated that successful students tend to spend more time on challenging concepts, indicating that persistence and focus are the key components of effective self-regulation. This aligns with Maldonado-Mahauad et al. (2018), who emphasize the importance of consistent planning, monitoring, and reflection in the context of self-paced MOOCs, further reinforcing the value of SRL behaviors in online learning success. On a broader level, Ergen and Kanadlı (2017) confirmed the large positive effect of SRL strategies on academic achievement across various disciplines and education levels, while Emagnaw (2019) cautioned about the importance of self-regulation and time management for high achievers. Collectively, these studies argue for supporting SRL through targeted feedback, as suggested by Butler and Winne (1995), and addressing challenges in real-time (Arakawa et al., 2021) can potentially enhance student outcomes.
2.3. At-risk student detection
Current attempts at identifying at-risk students in CS primarily rely on predictive learning analytics to detect low-performing individuals (Hellas et al., 2018). Such analytics are usually based on attempting to classify students as passing or dropout using some measurable features about the student (Prenkaj et al., 2020). Hellas et al. (2018) outlined various predictive features that has been used, including demographic, academic, and behavioral factors. For instance, Naseem et al. (2019) addressed high attrition rates in CS using 23 attributes encompassing students’ demographics, academic performance, financial information, and prior educational history. Quille and Bergin (2019) in turn demonstrated that student success in CS1 can be predicted with high accuracy using self-efficacy and mathematical ability metrics as key predictors. Further examining at-risk indicators, Ifenthaler and Yau (2020) identified crucial factors such as GPA, learning history, and clickstream data. Similarly, Ifenthaler et al. (2022) found that increased engagement with self-assessments correlates with better exam performance, underscoring the importance of integrating assessment data into performance predictions.
More recently, Sghir et al. (2022) noted the increasing use of machine and deep learning models in predicting academic outcomes, paving way for using more performance data in predictions. Waheed et al. (2023) demonstrated high effectiveness of deep learning techniques in predicting students at risk of failing in self-paced online education. Focusing on privacy concerns, Van Petegem et al. (2022) presented a privacy-friendly framework for early detection of students at risk of failing introductory programming courses. Utilizing interpretable machine learning techniques, they achieved high predictive accuracy early in the semester. A further study by Zhidkikh et al. (2024) reproduced these results using the same framework in a different CS1 course, highlighting the generalizability and explainability of at-risk student predictions through submission metadata and noting the importance of student self-evaluations in improving early dropout predictions. Finally, Moreno-Marcos et al. (2020) investigated dropout prediction in self-paced MOOCs incorporating self-regulated learning strategies as predictors. They found that event-based SRL strategies, derived from learners’ interactions with course materials, had high predictive power, outperforming traditional demographic and self-reported measures. Their result emphasizes the importance of considering complex, activity-based features over conventional data in predictive models.
Collectively, these studies underscore the critical role of predictive analytics in identifying at-risk students in CS education. They highlight the shift toward incorporating complex, activity-based features and advanced machine learning techniques to enhance predictive accuracy.
3. Research context and methods
The present study was conducted at the university of the first author which provides education and conducts research in the natural and social sciences. Students begin their studies with a major subject and have the ability to personalize their study plans by pick minor subject courses from other disciplines at the university. Computer Science (CS) is one of the subjects available as a major or a minor for the university’s students. First-year CS students usually complete the introductory CS1 and CS2 programming courses along with more theoretical courses on algorithms, data management, and data networks.
3.1. Course context
This study focused on the 6-credit (ECTS) CS1 course taught during fall 2023. The course is aimed primarily for first-year CS majors, but the course is open to all students at the university. In fall 2023, 321 students enrolled, with a pass rate of 63.5%. The 11-week course included multiple weekly activities, diverse learning resources, and a rigorous grading scheme. To pass, students were required to complete at least 40% of the weekly tasks, complete a course assignment (a mini-game project) and attend the final exam. The final grade was determined based on the number of completed weekly tasks and performance in the final exam. The course design allows getting the best grade even if either the final exam or weekly tasks were passed with minimal marks by compensating one with another.
Figure 1. Course week design of the studied CS1 course. The course design follows a weekly pattern, in which students first learn the week’s content, then apply the new content knowledge in homework, and finally, review their answers before proceeding to new content.
The course followed a weekly schedule, with a suggested learning strategy depicted in Figure 1. Each week began with two lectures for which the students could prepare by reviewing the week’s materials. The lectures were held in-person and online, during which the lecturer and students dynamically interacted through quizzes, interactive examples and discussions on the week’s topic.
Following the lectures, students completed a weekly task set consisting of 10 to 20 graded exercises of varying difficulty levels: introductory tasks (e.g., simulators and quizzes), basic programming tasks, bonus tasks (requiring more work), and guru tasks (requiring seeking extra knowledge beyond the week’s material). Students generally had the freedom to choose which exercises to submit, except for one or two “core tasks” that were mandatory each week. They had access to several learning resources, such as a coursebook, interactive examples, lecture recordings, and documentation. Students were encouraged to use only the course materials and not search the internet, as all necessary and verified materials are available within the course.
In addition, students could attend group study sessions in which they could study and get live help from teaching assistants. During the course, students were also required to participate in three one-on-one supervision sessions with a teaching assistant to assess the progress on the course assignment. Each course week ended with a task review session held after the task deadline, allowing students to compare each other’s answers. After the session, the students were encouraged to review and fix their answers using model answers.
3.2. The study
The study was implemented in two phases. The initial study entailed completing the course and using the available learning resources. During the study, the trace log data, task points, exam points, and final grades of the participants were collected from the virtual learning environment. After the course, an additional elective post-course questionnaire was administered to the original participants to gain additional self-report data. We followed the university’s guidelines for conducting studies involving human participants. Prior to participation, students were shown a full privacy statement on how their data were to be used in this study. Participating in both studies was fully elective, and participants could leave the study at any time. Participation was based on informed consent, and whether a student participated in the study carried no incentives or penalties. Altogether, $N=47$ students participated in the initial study. Out of them, 29 students also answered the additional post-course questionnaire.
3.2.1. Data collection
The main data of the study are the trace logs of the participating students. The initial logs were collected as raw action logs which contained timestamps of every action a student performed in the virtual learning environment during the 11 weeks of the course. The raw logs were processed into final trace logs by removing actions irrelevant to the course and by recoding the actions into context-relevant event codes. The codes relate to the various learning resources (e.g., course materials, lectures, supplementary materials, group sessions, one-on-one sessions) students use and the different tasks (introductory, basic, core, bonus, guru) students completed during CS1. Additionally, the events relating to different weekly task sets were differentiated, as student could, for instance, view and complete tasks from the current week or review and check answers to tasks from a previous week. Overall, 82 distinct event codes were used to describe student behavior during CS1. The description of all event codes and the specific approach for processing the raw logs is described in Appendix A.
In addition to trace log data, task completion and performance data was also collected for the duration of the whole course. The data included timestamps of all task submission attempts sent to the virtual learning environment and the correctness of each submission as determined by whether the student got full points from the task. Additionally, general performance data, such as weekly task points before the weekly deadline, task points after the deadline (as students could edit their answers based on the model answers), final exam points and final course grade were collected.
Finally, self-reports of students’ SRL were collected via a questionnaire after the course but before the final exam. The questionnaire included seven open-ended questions to which students could answer in brief sentences. Such an approach was selected to reduce the length of the questionnaire while allowing students to more broadly reflect on their SRL during CS1. The questions were modelled after MSLQ (Pintrich et al., 1991) which is a commonly used self-report tool for assessing SRL skills (Duncan and McKeachie, 2005). The full questionnaire is included in Appendix B.
3.2.2. Data analysis
Figure 2. The procedure for analyzing trace logs used in the present paper. Analysis follows Winne and Marzouk (2019)’s characterization of SRL, whereby student’s learning strategy are described as a transition of learning tactics. The approach for identifying learning tactics from trace logs is based on Matcha et al. (2019).
Two analyses were conducted to address the research questions. First, process mining used to examine students’ weekly learning strategies with course materials. Weekly dropout risk was calculated using a pre-trained predictor based on task completion data. Low- and high-risk strategies were then clustered to identify different learning types. Second, student profiles derived from these strategies were clustered and evaluated using self-reports for validation. Figure 2 outlines the full analysis procedure, with the details provided next. All anonymized data and analysis code is available online [^1] URL available upon request from the first author..
Before the main analysis, students’ weekly behavior patterns were extracted from trace logs. Building on the characterization of student SRL behavior via learning tactics and learning strategies (see Section 2.1), we first identified students’ learning tactics of using the course materials, and then we mapped strategies for applying the tactics during each course week. Since the course follows a weekly structure with consistent learning goals and tasks, we mapped strategies for each week. A process mining and clustering method, adapted from prior work (Zhidkikh et al., 2023), was employed. This method, based on event-based SRL analysis by Matcha et al. (2019) and Jovanović et al. (2017), involved splitting trace logs into per-student study sessions using the time distance between the events. Study sessions represent consecutive active interactions with course materials, such as student reading or completing a task. If the time difference between two events was longer than 25 minutes, the first event was set as an end of the previous session and the second as the start of a new session. We chose a 25-minute cut-off based on the number of weekly tasks and the average time students were expected to use on each task. Single-event sessions were then excluded to remove logs of accidental virtual learning environment interactions. Each session was represented as a frequency vector $(\text{fE}_{1},\text{fE}_{2},\ldots,\text{fE}_{82})$, where $\text{fE}_{i}$ indicates the relative frequency of an event type in the session. These vectors were clustered using $k$-medoids with the $\ell_{1}$ metric (Schubert and Lenssen, 2022) to enhance robustness of the analysis in case the dataset was too limited in size. The clusters were then interpreted as distinct learning tactics employed by students in CS1. Finally, each student’s weekly strategy was constructed by remapping the study sessions into tactic types and generating First-Order Markov Models of weekly tactic use through process discovery (Berti et al., 2023).
Additionally, the open-ended questionnaire answers were preprocessed for later analysis. The last two authors of this study individually coded initial answers into themes. The coding of the themes was not guided, but emerged from the questionnaire answers (cf. conventional content analysis in Hsieh and Shannon, 2005). Additionally, the two authors recorded whether a student had a self-reported positive, neutral or negative opinion of their learning behavior during the course. After individual coding, the two authors convened in a two-hour session to form a consensus on codings which were not initially agreed upon. The resulting codes of all students are thus jointly decided between the last two authors. The codes were used in subsequent analyses to compare students and characterize their behavior.
To answer RQ1 and RQ2, weekly learning strategies were paired with weekly dropout predictions to identify patterns in the behavior of low- and high-risk students. Since students may adjust their strategies weekly in response to environmental factors and prior performance (e.g., Pintrich, 2000), their dropout risk may fluctuate over time. We therefore assessed dropout risk weekly with a model developed in a prior study (Zhidkikh et al., 2024), which enabled the identification of “low dropout risk” and “high dropout risk” strategies. The used model was trained on task performance data from seven past iterations of the CS1 course, allowing prediction of weekly dropout based on task performance. Given the task performance data, the prediction model gives a dropout prediction $p_{\text{drop}}$ as a probability value from 0 (likely to pass) to 1 (likely to dropout). The learning strategies, modeled as First-Order Markov Models, were categorized into low and high dropout risk groups using the weekly predictions. These groups were then clustered using the Expectation Maximization algorithm. The cluster counts were selected by the first author using cluster validation indices (CVIs) as a guide. The first author then devised a dashboard, in which for each cluster median dropout probability, median grade of students, combined self-report codes, combined heuristic net (Weijters et al., 2006), per-strategy weekly performance data, per-strategy First-Order Markov Model, and per-strategy SHAP waterfall plot explaining the dropout prediction (Lundberg and Lee, 2017) were made available for the first two authors to help characterize the learning strategy types. To answer the research questions, the clusters were interpreted as strategy types by comparing them to each other using the aforementioned data. The interpretation was done by the first two authors through two hours of joint work, where the clusters were reviewed and the interpretations agreed upon.
Additionally, the strategy types and student self-reports were utilized to formulate students’ course activity profiles to further address RQ1 and RQ2, and also RQ3. First, weekly student strategies were mapped to the strategy types identified in the analysis of RQ1 and RQ2. Then, each student was assigned an activity profile represented as a frequency vector $(\text{fStrat}_{1},\text{fStrat}_{2},\ldots,\text{fStrat}_{N})$, where $\text{fStrat}_{i}$ represents the frequency of a student’s weekly strategies corresponding to a specific learning strategy type, and $N$ is the number of strategy types identified earlier. In other words, these profiles capture students’ course activities and their strategy adjustments. The use of frequencies simplified profiles, enhancing robustness of the analysis. Agglomerative clustering with the $\ell_{1}$ metric was applied to cluster students based on their activity profiles. The final cluster count was selected by the first author using CVIs and the resulting dendrogram. The resulting clusters were interpreted by the first two authors iteratively in up to three hours of joint work using activity profiles, mean dropout prediction, course performance, and coded self-reports, with the last data source particularly aiding cluster interpretation. To support comparison of performance data between clusters, Brunner-Munzel test for stochastic equality ($H_{0}:P(X<Y)=P(X>Y)$) was used as a robust non-parametric alternative to the Mann-Whitney U test, suitable for low number of observations and the presence of tied values (Neubert and Brunner, 2007). This paper adopts the reporting guidelines proposed by Karch (2021), including using stochastic superiority estimate $\hat{p}^{\prime\prime}$ for effect size. While the effect size estimates for Brunner-Munzel test do not have standardized significance categories, $\hat{p}^{\prime\prime}$ can be interpreted as the probability that random variable $X$ is lower than random variable $Y$ when splitting ties (i.e. observations of equal values) evenly (Karch, 2021).
4. Results
4.1. Trace log preprocessing and learning tactics
Figure 3. Cluster validation for clustering of student sessions. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better.
Code N Description Example codes in sessions
F_CourseMat_Examples 509 Focusing on course materials and examples in the materials
answer:materials-book-example
read:materials-book
F_Lec_Engaged 766 Focusing on the lecture and lecture materials
read:lecture-cur-wk
answer:lecture-cur-wk-example
F_Lec_Video 469 Focusing on lecture videos
read:lecture-cur-wk
watch-video:lecture-cur-wk
F_CurTasks_Intro 350 Focusing on current week’s introductory tasks
answer-wrong:tasks-cur-intro
answer:tasks-cur-intro
F_CurTasks_Core_Attempt 356 Focusing on current week’s core tasks, submitting incomplete solutions
answer-wrong:tasks-cur-core
read:tasks-cur
F_CurTasks_Core_Correct 410 Focusing on current week’s core tasks, submitting correct solutions
answer:tasks-cur-core
read:tasks-cur
F_CurTasks_Basic_Attempt 369 Focusing on current week’s basic tasks, submitting incomplete solutions
answer-wrong:tasks-cur-basic
read:tasks-cur
F_CurTasks_Basic_Correct 529 Focusing on current week’s basic tasks, submitting correct solutions
read:tasks-cur
answer:tasks-cur-basic
F_CurTasks_Extra 640 Focusing on current week’s extra tasks (bonus, guru) or supplementary tasks
read:materials-book
read:tasks-extra
TA_Sess 57 Attending one-to-one session with TAs
session-start
session-end
F_PrevTasks_Basic 361 Focusing on reviewing last week’s tasks, focusing on fixing basic answers
read:tasks-prev
answer:tasks-prev-basic
answer-wrong:tasks-prev-basic
F_PrevTasks_Deep 629 Focusing on reviewing last week’s tasks, reviewing and fixing multiple tasks
read:tasks-prev
answer:tasks-prev-core
answer:tasks-prev-intro
Table 1. Weekly learning tactics interpreted from student session clusters. For each tactic, short code, number of exemplary student sessions and description of the tactic is given.
Overall, the trace logs were split into 7,226 individual student sessions. After removing sessions with just one event, 5,445 sessions were included in the analysis. To select a reasonable number of session clusters, multiple cluster validation indices (CVI) were computed. The clustering performance of the $k$-medoids clustering on the student sessions is depicted in Figure 3. Given the weekly schedule of the course (Figure 1), we expected to select at least 7 session types. We ultimately selected $K_{\text{sess}}=12$ session clusters, as the CVIs for this cluster count are close to best and because a larger number of clusters would allow capturing more detail in students’ weekly strategies. From the session clusters, we interpreted 12 tactics the participants used during CS1. The descriptions of all tactics, the number of student sessions and some example event codes in the student sessions are listed in Table 1.
It must be noted, that while the table shows distinct event codes for each tactic, in practice student sessions may contain a mix of similar events. For example, both “Focusing on current week’s basic tasks, submitting correct solutions” and “Focusing on current week’s basic tasks, submitting incomplete solutions” tactics have examples of students answering wrongly (answer-wrong:tasks-cur-basic) and correctly (answer:tasks-cur-basic), but the proportion of the correct answers in the sessions is much higher for the former tactic (32% answering correct vs. 7% answering wrong) than the latter tactic (10% answering correct vs. 37% answering wrong). Therefore, the relative proportions of recorded event types within each cluster were primarily used to interpret the student session clusters as tactics.
4.2. Weekly learning strategy types
Figure 4. Cluster validation for clustering of the weekly learning strategies. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better.
Figure 5. The identified weekly strategy clusters; each cluster shows multiple weekly learning strategies as vertical bars; colors in weekly learning strategies indicate learning tactics listed in the figure legend and described in more detail in Table 1; the size of each area describes the relative frequency of the used tactic within the strategy; for each cluster, interpreted strategy type, number of strategies in the cluster $N_{\text{strat}}$ and average predicted dropout probability $\bar{p}_{\text{drop}}$ is given.
(a) Heuristic Net of strategies in the cluster that was identified as Falling behind, realizing struggle too late.
(b) SHAP waterfall plot explaining prediction for a weekly learning strategy of a single student inside the cluster that was identified as Falling behind, realizing struggle too late.
Figure 6. Example of graphs used to identify learning strategy type Falling behind, realizing struggle too late.
After recoding student sessions using the learning tactics and splitting sessions by course week, we ended up with 458 weekly learning strategies for further analysis. After applying weekly dropout predictions to group weekly learning strategies, 396 strategies were categorized as low dropout risk strategies ($p_{\text{drop}}\leq 0.5$) and 62 as high dropout risk strategies ($p_{\text{drop}}>0.5$). Initial cluster validation for the Expectation Maximization algorithm using BIC suggested using a diagonal covariance matrix. Based on further evaluation with multiple CVIs (Figure 4), we selected $K_{\text{pass}}=3$ clusters for low dropout risk strategies and $K_{\text{drop}}=9$ clusters for high dropout risk strategies. Here, we preferred to select a larger cluster number to better capture various learning strategy types. To identify the learning strategy types, we plotted the frequencies of the learning tactics used in every weekly strategy, resulting in Figure 5. In addition, we generated heuristic nets for each cluster to aid interpretation. In some edge cases, we also reviewed the predictions of the used dropout prediction model using SHAP waterfall plots. Examples of both are provided in Figure 6. Overall, we interpreted the following learning strategy types from clusters of low dropout risk strategies:
  1. (1)
    Task-oriented, focused on performance: Student invests significant time in completing mandatory and basic tasks, ensuring the main weekly goal is finished.
  2. (2)
    Seeking understanding: Student prioritizes gaining understanding through extensive use of lecture materials and supplementary resources before and during task completion.
  3. (3)
    Resource-focused, more time on materials than tasks: Student focuses first on course materials rather than task completion, consistently attending lectures and reviewing model answers.
In turn, we interpreted the learning strategy types below from clusters of high dropout risk strategies:
  1. (4)
    Falling behind, realizing struggle too late: Student views previous weeks’ model answers without engaging in current tasks. The median grade of students inside the cluster is 0, indicating that this strategy type is used by students near dropout.
  2. (5)
    Attempting with examples: Student relies on examples to attempt tasks, possibly experimenting with different tactics to find a suitable approach.
  3. (6)
    Low engagement: Student relies solely on lectures as a learning resource, completing few tasks with points below the weekly goal. No effort used to seek help from other course materials.
  4. (7)
    Late reliance on model answers: Student depends on viewing model answers from previous weeks retroactively to catch up with the course pace, with minimal engagement with other materials.
  5. (8)
    Slow start, finding study pace: Student initially completes few random tasks despite using a variety of materials. Occurred only in the first week, representing an initial struggle to find the study rhythm.
  6. (9)
    Struggling to understand, needing help: Student spends considerable time searching through course materials in an effort to understand (cf. Seeking understanding) but fails to grasp the concepts, completing fewer tasks.
  7. (10)
    Superficial review, stuck on a single resource: Student engages primarily in watching videos without utilizing other support materials or progressing to task completion.
  8. (11)
    Browsing materials with sporadic task attempts: Student focuses on supplementary materials and tasks as the main activity, making random attempts at tasks, possibly to grasp weekly content but generally completing few or no actual tasks.
  9. (12)
    Risky focus on mandatory tasks only: Student concentrates solely on mandatory tasks with minimal involvement in other course activities, either only attending one-on-one supervision sessions or engaging very little otherwise, representing a risky approach to course completion.
4.3. Student learning profiles
(a) Cluster validation indices for student profile clusters. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better.
(b) Dendrogram of the clustered student profiles. The vertical axis includes identification numbers of each student. The vertical line indicates the selected cluster count $K_{\text{student}}=5$.
Figure 7. Cluster validation for student profile clusters using CVIs and a visual approach using a dendrogram.
Figure 8. Student profile clusters displayed as the number of students that used each weekly strategy type throughout all course weeks. For each cluster, the interpreted profile type name, number of students in the cluster $N_{\text{student}}$ median percentage of the completed weekly tasks $\text{Md}(w)$, median percentage of exam points $\text{Md}(w)$, median course grade on scale 0–5 $\text{Md}(g)$ and mean dropout prediction over all weekly strategies $\bar{p}_{\text{drop}}$, is listed. For clusters with students who filled out the self-report, themes mentioned by at least half of the answerers are listed (see Appendix C for a list of all themes and descriptions). Additionally, a distribution of self-reported opinions of own effort are shown. Cluster interpreted as Course dropouts does not include any students with self-reports.
We used CVIs and a dendrogram to select the linkage criterion for clustering student profiles. Initial validation suggested to use complete linkage criterion. In further validation for cluster numbers, multiple CVIs (Figure 7(a)) suggested 4–6 student profile clusters. Based on a visual review of a resulting dendrogram (Figure 7(b)), one student’s profile (Student 1) appears to strongly differ from the other students. As such, we ultimately selected $K_{\text{student}}=5$ as the cluster number, with four clusters representing learning profile types and one cluster containing the outlier student.
An overview of the interpreted learning profile types along with a summary of the available self-report data is shown in Figure 8. A further listing and description of all emerged themes used in the study’s analysis is listed in Appendix C. Overall, three of the learning profiles represent various levels of activity during CS1, while one learning profile includes all students that dropped out mid-course. Further, the profile of the single outlier student represents a case of a low-activity student who managed to pass the course by compensating low task performance with good exam results. All clusters include students without self-reports, with the cluster of course dropouts having no recorded self-reports at all. The profile of a single student contains a self-report that was extensively used in interpreting the profile. Nevertheless, visually, there appear to be some differences in self-reported opinions of students’ own performance between the clusters. Next, we describe the learning profile types interpreted from the clusters based on the visualizations, descriptive statistics, confirmatory tests, and excerpts from the self-reports.
4.3.1. Adaptive understanding seekers
The first interpreted profile type ($N_{\text{student}}=20$) contains students who primarily sought understanding and adapted their study strategies as needed. Generally these students have stuck to the understanding-seeking strategy throughout the course. However, some students do appear to have varied their strategies during the course to adjust to new topics or find other ways of learning, as exemplified in the quotes below:
[Student 10]: [My] routines for setting the week’s rhythm improved during the first course week. Initially, I did double the work a bit too much, e.g., re-watching lecture recordings multiple times because I did not read the course materials first.
[Student 33]: As the course progressed, I read the materials less and less. I started reading theory texts only when completing the weekly tasks. I felt I learned the best by applying the written information right away.
Some students occasionally slipped into high dropout risk strategies, but eventually managed to recover. Based on the self-reports, many students in this profile type used various resources, including external sources from the internet while remaining critical. An example is provided via a quote below:
[Student 3]: I survived [the course] by mainly using the course materials and attending weekly group sessions. Also, the peer support group motivated and “tied” me to completing the weekly tasks. I peeked at some info via a search engine; I hope was able to differentiate between valid and invalid guides.
As such, the profile represents students who really wanted to ensure understanding of the topic while progressing through the course. It must be noted that not all understanding seekers succeeded with high grades; those who self-reported bad experiences with the course attributed it to external factors such as course workload and personal life:
[Student 19]: The course has a high workload and for me required much more effort than 10–12 hours weekly.
[Student 12]: When doing the course while employed, you can’t use all your time for the course. On the other hand, studying is a good counterbalance to work.
In addition, two students (Student 8 and Student 28) did not complete all course requirements in time, which prevented them from getting the final grade despite being otherwise active during the course. Student 28 filled a self-report, where they explicated the struggles as follows:
[Student 28]: This fall was very busy, and during some weeks I did not make as much effort as needed to deeply learn the topic.
[…]
4.3.2. Diligent course content followers
The second profile type ($N_{\text{student}}=14$) represents students who conscientiously used the course materials and appeared to closely mimic the suggested weekly schedule (cf. Figure 1). Their primary strategy was material-focused during the majority of the course duration. However, toward the end of the course, more students shifted to the understanding-seeking strategy. This possibly indicates some students’ need to enhance comprehension before final exams as the example course schedule may not have been enough to gain sufficient understanding of the topics. In self-reports, students emphasized the importance of seeking help:
[Student 2]: I read course materials, I made a learning map, I sought help via Google and asked for help from the people I know.
[Student 5]: …my partner who knows how to code helped me a lot.
[Student 23]: I was in a peer support group, which helped me to schedule the weekly tasks. I also asked for help from friends a lot.
Students also reported relying heavily on the teaching assistants, implying some dependence to complete tasks and even to self-regulate:
[Student 40]: I used over 20 hours per week on the course but didn’t understand anything without group sessions and teaching assistants. […] I don’t even dare to attend CS2 this spring since I’m so mentally exhausted.
[Student 43]: I felt I was left by myself with my problems the first time studying CS1, so now I tried to make use of all the available help channels.
Overall, while following the ideal schedule worked for some students, others reported the course to be challenging and struggled to adapt their strategies, resulting in varied outcomes.
[…]
4.3.3. Active task-focused learners
The active task-focused learner profile type ($N_{\text{student}}=5$) represents students who focused primarily on completing the weekly tasks. Unlike the two aforementioned profiles, these students consistently employed effective, low dropout risk strategies throughout the course, with only one student having initial adjustment struggles during the first week. The active focus on course tasks can be seen in low dropout predictions of weekly strategies and the high median percentage of completed weekly tasks. The active task-centered approach is reflected in the self-reports, exemplified by the following quote:
[Student 36]: At the start of the week I read the coursebook, studied the weekly tasks and completed them since I knew how to do them.
However, high task activity does not guarantee a high grade. One student mentions language barrier as an external obstacle for their performance:
[Student 18]: This was my first year studying CS. Because of this and the language barrier, I had a lot of problems. Because of this, self-study worked better for me. I did not find help from group supervision sessions.
[…]
4.3.4. Persevering until the end
A notable outlier to all students is Student 1, who completed the course with the best grade despite only minimally interacting with weekly tasks. Even when relying only on high dropout risk strategies, the student achieved the best grade by obtaining near best points from the final exam. While the final grade can be seen an indicator of good performance, the student’s self-evaluation of own learning methods is much more critical, with them setting a low bar for own performance:
[Student 1]: [The study style] was not really working. The worst for me was time management, which required extra effort. For now, my goal is to just pass the course for which my approach should be good enough.
When further reflecting on regulating own effort, the student brought up a larger issue with transitioning to university:
[Student 1]: …I had difficulties getting accustomed to university studies. At the same time I had to also get used to living by myself.
Visually, the student’s strategy was generally focused on a quick and risky attempt to complete mandatory tasks, and in course week 9 there is are no observes student sessions at all. The risky approach to studying is confirmed in the self-report:
[Student 1]: I usually used one day for programming stuff. With the course assignment I used a bit more time in places.
When answering the question about scheduling learning, the student notes their reactive nature to deadlines:
[Student 1]: Actually, I went by deadlines and personal life, depending on how I felt or had time.
Further, the student studied alone, highlighting avoidance of social events as the main reason:
[Student 1]: I mostly studied remotely, to which I am not accustomed to, but large lecture halls don’t really seem attractive to me.
Nevertheless, the main reason to passing the course could be attributed to the student’s ability to find and information necessary:
[Student 1]: I mainly used course materials and the internet. I was doing pretty well in terms of finding the right information. […] The course materials were good.
4.3.5. Course dropouts
The final profile type ($N_{\text{student}}=7$) includes all participants who dropped out during the course. While no students in the profile filled a self-report questionnaire, the Brunner-Munzel tests confirm that both grades ($\hat{p}^{*}(39)=27.221,p<.001,\hat{p}^{\prime\prime}=0.975$) and dropout predictions ($\hat{p}^{*}(20.362)=14.408,p<.001,\hat{p}^{\prime\prime}=0.085$) of the students in the profile in are not stochastically equal to the other students. Evaluating the effect size, the students in this profile tend to have lower grades and higher predicted dropout than the other students.
Upon a manual review of all the students, we identified three paths to dropout:
  • four students dropped out during the first course weeks with only high dropout risk strategies used;
  • two managed to hang on the course until the course difficulty considerably increased in weeks 3 and 5;
  • one student started with low dropout risk strategies and managed to continue through initial complex topics, however they quickly fell to using high dropout risk strategies and dropped out right after a particularly difficult topic was introduced.
In other words, even if a student in this profile did manage to use low dropout risk strategies, ultimately every student reverted to high-risk strategies 1–2 weeks prior to dropout. In this profile, all students eventually dropped out by the midway of the course when programming tasks became more complex.
5. Discussion
This study aimed to examine the behavioral patterns of CS1 students throughout the course by analyzing trace logs and student dropout predictions, interpreting the results through the lens of SRL. The findings demonstrate that combining event-based logs, dropout predictions, and basic self-reports reveals clear differences between high and low dropout risk CS1 student behaviors, providing detailed insights into their successes and struggles throughout the course. Although detecting and understanding learning patterns associated with different dropout risks can help address course attrition, research into automatic pattern detection remains scarce in CSEd (Omer et al., 2023). This study therefore contributes to the research area by demonstrating a method for semi-automated learning pattern analysis that complements and extends qualitative inquiries into learning patterns of programming students, such as those by Liao et al. (2019). Furthermore, by relating trace log data to self-reports to obtain a more comprehensive narrative of student behavior during CS1, this study contributes to the discussions regarding practical applications of event-based measures (e.g., Moreno-Marcos et al., 2020; Fan et al., 2023; Han, 2023; Choi et al., 2023).
5.1. Addressing the research questions
Overall, we identified and interpreted 12 distinct weekly learning strategy types of various dropout risk (see Section 4.2) and categorized student course behaviors into five learning profile types (see Section 4.3). The results bear resemblance to those of Liao et al. (2019) and Moreno-Marcos et al. (2020), confirming the validity of the study. Further, the weekly learning strategy types relate to other studies in which SRL learning strategies were mapped in CS1 context (Falkner et al., 2014; Silva et al., 2024; Garcia et al., 2018).
Regarding RQ1, we identified three clusters among strategies with low dropout probability and interpreted them as the following weekly learning strategy types: task-oriented, prioritizing main task performance, seeking understanding, and resource-focused, prioritizing course materials over tasks. These strategies align with Liao et al. (2019), who reported that high-performing students tend to use more varied course resources. Furthermore, information seeking and experimentation with simpler tasks before solving more complex problems are common SRL strategies in computer science education, as exemplified by Falkner et al. (2014) and Silva et al. (2024). Although prioritizing task completion is not generally reported as an SRL strategy itself, it can be considered a form of goal-setting and planning, as students allocate more time and effort to tasks (e.g., Garcia et al., 2018). Additionally, from clustering student behaviors over the entire course, we identified three clusters that we interpreted as learning profile types associated with course success: adaptive understanding seekers, diligent course content followers, and active task-focused learners. These profile types relate to each other in that most of the students of these types applied low dropout risk strategies throughout the course, and often actively varying strategies depending on the weekly topic. This aligns with Liao et al. (2019), who noted that high-performing students tend to vary their study approaches depending on the task.
Regarding RQ2, we identified nine distinct clusters among strategies with high dropout probability from which we interpreted high dropout risk strategy types. Unlike the low dropout risk strategy types, these strategy types show greater variation in behavior and dropout risk. For instance, one of the interpreted strategy types, slow start, finding study pace, does not by itself indicate dropout, as study participants who employed this strategy improved their learning style as the course progressed. This strategy type aligns with the common observation that students may initially struggle to self-regulate in a new context but learn it over time (e.g., Loksa and Ko, 2016; Winne and Marzouk, 2019). Most of the interpreted strategy types, such as attempting with examples; low engagement; late reliance on model answers; and struggling to understand, needing help, represent in our view basic temporary unsuccessful strategies that even successful students may experience and recover from, as observed in the identified student profile clusters (cf. Figure 5). Such notion has been noted as well by Liao et al. (2019), who noted that lower-performing students can overcome problems just as high-performing students may temporarily struggle. However, the final interpreted strategy types — falling behind, realizing struggle too late; superficial review, stuck on a single resource; browsing materials with sporadic task attempts; and risky focus on mandatory tasks only — seem to describe behaviors of students on the verge of dropping out. In the case of the course dropouts present in this study, these strategy types were among the last ones those students attempted before dropping out. In our view, these strategy types can be related to issues with time management and difficulty assessment, both identified by Falkner et al. (2014) as the most frequent unsuccessful SRL strategies in CS. On the other hand, these strategy types can be simply seen as acts of giving up prior to dropout. In other words, those students who persevere through problems recovered from high-risk strategies, while dropouts remained locked in high-risk behavior until they were overwhelmed by the workload — a common narrative for dropping CS1 (Petersen et al., 2016; Kinnunen and Malmi, 2006).
Regarding RQ3, we found that SRL as we observed and interpreted via trace logs and students’ self-reported evaluations of their own SRL are generally consistent and complementary. Although we initially mapped learning strategies and student profiles using only trace logs and dropout prediction, self-reports enhanced the interpretation of the types. Specifically, both trace log-based data and self-reports were equally valuable but provided different perspectives. While Liao et al. (2019) and our self-reports offer a snapshot of students’ cognition at the end of the course, trace logs allow following the full learning path of a student and capture their overall learning profile as transitions between learning strategies (cf. Figure 5). Although the somewhat open-ended self-report tool we used may not be appropriate for predicting dropout on its own, it aids describing student behavior. In this study, the self-reports were instrumental in being able to differentiate the three main learning profile types: adaptive understanding seekers, diligent course content followers, and active task-focused learners. Furthermore, the profile of a student who passed the course with full marks despite being identified as high-risk (see Section 4.3.4) would not have a reasonable explanation without the student’s self-report. Thus, our results confirm an argument made in an earlier study (Zhidkikh et al., 2023): using predictive learning analytics and trace logs alongside self-reports allows for a more comprehensive understanding of how and why a student uses the available learning resources. Aligning with numerous prior studies (Tempelaar et al., 2020; Zhidkikh et al., 2024; Fan et al., 2023; Lim et al., 2021), we posit that the inclusion of self-reports can improve the interpretation of purely trace log-based analyses by contextualizing them and bringing in cognitive aspects that cannot be easily captured as atomic events.
5.2. Practical implications
We believe the results demonstrate students’ study habits and progress in a way that can be useful for educating TAs. This knowledge can, of course, be shared among teachers, while TAs are recruited to various roles in which they interact with students, including autonomous positions and providing guidance on the coursework (Mirza et al., 2019). In using the results as educative material, specific attention could be given to illustration of dropouts, as these cases show the point in the course over which students with unproductive strategies hardly progress and encourage early interventions. On the other hand, the outlier student profile (Persevering until the end) demonstrates that, regardless of strategies that appear risky, a student may achieve a good result and is hence likely to have good potential for finding and grasping the material (see illustration at the end of Section 4.3.4). From this perspective, a view that the course structure “upholds” beginner students (through allowing a focus on mandatory exercises only or late rework with them) may pay off for some students. In effect, the student in the outlier profile acknowledged difficulties in transitioning to university, illustrating that student difficulties need to be screened and addressed in a broader scope than strategies with learning to program.
An intervention between a TA and a student could begin gently, with the TA asking the student to reflect on their study habits in relation to the learning profile types (clusters in Figure 8). Beginning from the student’s reflection conforms to good dialogue-based counseling practice (Peavy, 1998), with TAs avoiding the situation of needing to be able to give “best” advice to students. The goal would be to prompt the student to consider whether their current strategies seem risky or successful. When necessary, the profiles identified as successful can be offered as suggestions for improving learning strategies. Such a use of the results would also serve to validate them, revealing to what extent the current profiles are applicable in practice or if more varied profiles exist among student cohorts.
We would also pay attention to students within the learning profile types generally considered to reflect successful strategies. We observed that they may yet achieve rather different grades. We anticipate that this condition derives from students’ varied backgrounds, which also calls for sensitive dialogue in supervision sessions in the scope broader than navigation through course exercises.
Learning analytics in programming courses have been noted to lack granularity, as efforts are often based on limited data, such as assessment results (Omer et al., 2023). The present study revealed paths of students progress through clustered profiles that became available by employing multiple types of data, including trace log data and weekly prediction data, along with the students’ end-of-course reflections. Above, we described one scenario of using these results as an informative source for supervision during subsequent courses. A future goal that naturally intrigues us as teachers is the possibility of seeing timely, automated characterizations of students’ current progress. This leads to the well-acknowledged challenges of how teachers can effectively use learning analytics data and how effective automated feedback can be generated (Shibani et al., 2020). We plan to continue by replicating the present study with larger samples and varied contexts, and then explore methods to automatically map new datasets to previously developed clusters. Shibani et al. (2020) noted the need to inform users of learning analytics outcomes about potentially misleading information. If the suggested future direction turns out fruitful, the automated results could be used safely in dialogue-based supervision sessions, where the nature of such analytics can be critically reflected on.
5.3. Limitations and threats to validity
We recognize several threats to validity in our study design. In this subsection, we have argued for our chosen research setting, its limitations, and how we have strived to mitigate potential threats to validity.
Relatively small number of participants: The number of participants in this study is relatively small considering the total number of students in the course, and considering that the data were analyzed with statistical methods. However, despite the number of participants in the statistical analyses (N = 47), the number of observations collected from 11 course weeks is higher, i.e., 5,445 student sessions grouped into 458 weekly learning strategies. Furthermore, while we used the Brunner-Munzel test mainly because our data failed to meet the assumption for independence of observations due to our study design, the Brunner-Munzel test has been shown more robust with smaller sample sizes than other adjacent non-parametric tests. Finally, SRL trace log analysis has been applied successfully on a relatively low number of participants before (Zhidkikh et al., 2023).
Changes in study profile: We categorized weekly learning strategies in our analyses given the weekly schedule of the course (see Section 3.1), yet this level of abstraction might not capture all nuances of changes in a student’s study profile. That is, changes in learning strategy are likely to change more often than on a weekly basis. This choice was by design to abstract the results to be more easily interpreted, yet the approach might overgeneralize the phenomenon of study profiles.
Arbitrary number of clusters: Arguably, the data, as visualized in Figure 5, could be clustered with a higher or lower number of clusters, the former increasing the precision on which the results are reported, and the latter abstracting student behavior. We chose the number of clusters based on the observations of the first author, who is deeply familiar with the design of the course in this study. Based on these observations, a different numbers of clusters were initially considered, yet 12 clusters captured detailed enough behavior while providing a sufficient level of abstraction for the results to be relatively easy to interpret.
Arbitrary study session length: We chose an arbitrary length of 25 minutes for a study session in the trace log analysis, which potentially skews the results in unforeseeable ways. We chose this number based on our previous observations on how much time it is expected for a student to complete one exercise. Some other studies (e.g., Jovanović et al., 2017; Matcha et al., 2019) have used a 30 minute cutoff, which is also arbitrary. Other heuristics could have potentially been more accurate, but there are no standardized rules for determining study sessions from trace logs (Kovanovic et al., 2015). Reliably determining when a study session ends would have required a more controlled experiment (e.g., a laboratory setting, or capturing participant screens). We hypothesized that such a requirement would have yielded considerably less participants.
Self-report answer interpretation: In addition to the statistical tests, we analyzed the self-reports of 25 participants using content analysis. While the results yielded by content analysis are subject to researcher interpretation and subjective bias, we strived to mitigate this by having two authors analyze the data independently, discussing possible discrepancies, and forming a consensus, as detailed in Section 3.2.2. Despite this approach, another group of researchers might have formed at least slightly different set of themes.
Single-cohort study: The participants were recruited from a single cohort in a single university course. This approach poses potential challenges to the generalizability of the results concerning, e.g., topics besides CS1 or computing in general, cross-cultural differences, different teaching approaches, or some intra-cohort intricacies not considered in this study. As with many other research settings, these potential differences may be considered in replication studies.
6. Conclusion
In this study, we investigated the behavioral patterns of students in a CS1 course and their interactions with the course’s learning resources. By applying trace log analysis and student dropout prediction methods developed in two prior studies (Zhidkikh et al., 2023, 2024) to a cohort of CS1 students, we identified and interpreted 12 weekly learning strategy types based on students’ interactions with course resources and four core learning profile types that describe how these interactions change over the course duration. Among these learning strategy types, three were interpreted from low dropout risk behaviors and nine from high risk behaviors, indicating that dropout risk can be observed in various forms and may require different types of interventions. We compared the interpreted strategy types with other studies on self-regulated learning, finding close connections to related works such as Liao et al. (2019). Additionally, our observations of student behavior prior to dropout was consistent with prior studies that map how students drop CS1 (e.g., Kinnunen and Malmi, 2006; Petersen et al., 2016). Altogether, previous studies provide initial support and validation to our interpretations.
Further, we found points of contact between student learning profiles and self-reports, validating our interpretations of trace log data. Overall, the combination of trace logs and dropout predictions enhanced the identification and interpretability of successful and at-risk student behaviors. Given that dropout prediction can be performed multiple times throughout the course, this study demonstrates a semi-automated method for obtaining detailed narratives of study behavior and associated dropout risk during the course. Notably, this research approach is easily applicable to large CS1 course implementations where trace logs and basic task performance data are available. Through this approach, educators can better understand student study behaviors to devise more timely and targeted interventions.
Nevertheless, this study has limitations that warrant further studies. Most importantly, different CS1 contexts and with larger cohorts is needed. Given that CS1 course we obtained participants from had 321 enrolled students, there may be more unique learning paths that our identified strategy types and learning profile types do not fully encompass. Different CS1 course contexts may reveal additional nuances in student behavior, particularly concerning dropout. Additionally, we see potential for a reproduction study (cf. Ihantola et al., 2015, , Figure 2) by examining paths to dropout using longitudinal analysis tools. For example, applying Latent Transition Analysis (Lanza et al., 2003) to study sessions could uncover more intricate learning paths leading to dropout.
Appendix A Trace event codes and the detailed coding approach
Event code Description
answer-wrong:lecture-cur-wk-example The student answered wrong to an example question in the current week’s lecture materials.
answer-wrong:materials-book-example The student answered wrong to an example question in the book materials.
answer-wrong:tasks-[prev|cur]-self-assessment The student answered wrong to a self-assessment task in the previous/current week.
answer-wrong:tasks-[prev|cur]-intro The student answered wrong to an introductory task in the previous/current week.
answer-wrong:tasks-[prev|cur]-basic The student answered wrong to a basic task in the previous/current week.
answer-wrong:tasks-[prev|cur]-core The student answered wrong to a core task in the previous/current week.
answer-wrong:tasks-[prev|cur]-bonus The student answered wrong to a bonus task in the previous/current week.
answer-wrong:tasks-[prev|cur]-guru The student answered wrong to a guru task in the previous/current week.
answer-wrong:tasks-[prev|cur|extra|supp]-supplementary The student answered wrong to a supplementary task in the previous/current week or in extra/supplementary materials.
answer:lecture-[prev|cur|next]-wk-example The student answered to an example question in the previous/current/next week’s lecture materials.
answer:materials-book-example The student answered to an example question in the book materials.
answer:tasks-[prev|cur|next]-self-assessment The student answered to a self-assessment task in the previous/current/next week.
answer:tasks-[prev|cur|next]-intro The student answered correctly to an introductory task in the previous/current/next week.
answer:tasks-[prev|cur|next]-basic The student answered correctly to a basic task in the previous/current/next week.
answer:tasks-[prev|cur|next]-core The student answered correctly to a core task in the previous/current/next week.
answer:tasks-[prev|cur|next]-bonus The student answered correctly to a bonus task in the previous/current/next week.
answer:tasks-[prev|cur|next]-guru The student answered correctly to a guru task in the previous/current/next week.
answer:tasks-[prev|cur|next|extra|supp]-supplementary The student answered correctly to a supplementary task in the previous/current/next week or in extra/supplementary materials.
check-model-answer:tasks-[prev|cur]-basic The student checked the model answer for a basic task in the previous/current week.
check-model-answer:tasks-[prev|cur]-core The student checked the model answer for a core task in the previous/current week.
check-model-answer:tasks-prev-bonus The student checked the model answer for a bonus task in the previous week.
check-model-answer:tasks-prev-guru The student checked the model answer for a guru task in the previous week.
check-model-answer:tasks-prev-intro The student checked the model answer for an introductory task in the previous week.
check-model-answer:tasks-prev-supplementary The student checked the model answer for a supplementary task in the previous week.
[join|leave]-lecture:lecture-cur-wk The student joined/left the current week’s lecture.
read:lecture-all The student read general lecture materials.
read:lecuture-cur The student read the lecture materials during the lecture.
read:lecture-[prev|cur|next]-wk The student read the previous/current/next week’s lecture materials.
read:materials-book The student read the book materials.
read:tasks The student read general task materials.
read:tasks-cur The student read the current week’s task materials.
read:tasks-extra The student read extra task materials.
read:tasks-next The student read the next week’s task materials.
read:tasks-prev The student read the previous week’s task materials.
read:tasks-pre The student read pre-course task materials.
read:tasks-review The student read task review materials.
read:tasks-supp The student read supplementary task materials.
session-start:None The student started a TA one-on-one session.
session-end:None The student ended a TA one-on-one session.
watch-video:lecture-[prev|cur]-wk The student watched the previous/current week’s lecture video.
watch-video:tasks-review-[prev|cur|next] The student watched the previous/current/next week’s task review video.
workshop-[start|end]:None The student joined/left a group study session.
Table 2. All event codes used in the trace logs analyzed in this study. The table is condensed by combining variations of the same code into the same row.
The trace logs used in the study were generated from raw logs obtained from the virtual learning environment. The original logs contained rows of newline-separated entries in the following format:
    2023-09-01 12:00:00,000 INFO: username [127.0.0.1]: ACTION
where ACTION was some action related to a students’ interaction with the virtual learning environment. The original logs were already limited to only the study participants and only to the duration of the CS1 course.
Original logs were further filtered and re-coded into trace logs as follows:
  1. (1)
    If the action was not related to the studied course (e.g., not a course task, not a course document, general interaction with the system), the action was discarded.
  2. (2)
    Actions were re-coded into initial event codes. Course week information was added to some codes (e.g., “a student views tasks from week 1” and “a student views tasks from week 2” were separate events).
  3. (3)
    Events with week information were simplified into “current week” and “previous week” events (e.g., “a student views tasks from week 2” was simplified “a student views current week’s tasks” if the event occurred during course week 2).
  4. (4)
    Some consecutive events of the same type were merged. For example events of type “a student views tasks” were merged since the event would occur every time a student scrolls the task page.
  5. (5)
    The final trace logs were generated in format TIMESTAMP;STUDENT;EVENT_CODE.
The final used event codes are listed in Table 2.
Appendix B The post-course questionnaire
[Note: Text within brackets are Authors’ comments.]
Reflect briefly on your learning style and on the course in general. Answer the questions below with at least two sentences per question. Please especially reflect on what kinds of course materials you tried and whether they were beneficial in studying the course.
  1. (1)
    What was your usual study week in this course like? [Probing for general learning strategies.]
  2. (2)
    How did you schedule your studies during the course weeks? [Probing for time management.]
  3. (3)
    How did you feel about your ability to work towards the course? [Probing for effort regulation.]
  4. (4)
    What support channels and materials did you use? [Probing for peer learning & help-seeking strategies.]
  5. (5)
    Did you change your study methods during the course? Why? How? [Probing for metacognitive self-regulation.]
  6. (6)
    How effective did you find your study methods? [Probing self-evaluation.]
  7. (7)
    Do you have any other thoughts on studying in the CS1? [Probing extra insights.]
Appendix C Themes identified from questionnaires
Theme $N$ Description
Course Materials 24 The student indicates using the provided course materials.
Weekly Assignments 23 The student reports completing weekly assignments.
Lecture Participation 21 The student reports attending lectures or viewing lecture materials.
Emphasis on Regularity 20 The student highlights the importance of regular study habits.
Group Sessions 19 The student reports attending computer lab sessions.
Positive Opinion of Own Effort 16 The student gives an overall positive opinion of their own effort in the course.
Peer Support Groups 15 The student reports participating in a guided peer support group during the course.
Additional Online Materials 15 The student reports utilizing supplementary online resources.
Importance of Acquaintances 11 The student emphasizes the role of acquaintances in completing the course.
Anticipation of Future Challenges 11 The student exhibits self-regulation by preparing for upcoming topics or difficulties.
Perceived Course Workload 11 The student mentions that the course is labor-intensive.
Adaptation of Study Methods 10 The student adapts study methods to suit their individual needs.
Neutral Opinion of Own Effort 8 The student gives an overall neutral opinion of their own effort in the course.
Work-Induced Study Rhythm 6 The student indicates that employment provides structure to their studies.
Seeking External Help 5 The student reports seeking assistance from others to complete assignments.
Negative Opinion of Own Effort 5 The student gives an overall negative opinion of their own effort in the course.
Constraints from Other Courses 5 The student discusses how other courses limit their ability to focus on this one.
Externalizing Self-Regulation 4 The student relies on others (e.g., teaching assistants) for guidance and regulation.
Personal Life Constraints 4 The student brings up additional personal life factors that impose constraints.
Emphasis on Group Work Benefits 4 The student underscores the significance of group work in course completion.
Irregular Study Habits 3 The student mentions irregularity in their study activities.
Synchronization of Lectures and Assignments 3 The student emphasizes the importance of timing between lectures and assignments.
Challenges in Studying 3 The student brings up challenges in studying the course material.
Deviation from Established Practices 3 The student indicates deviating from their usual study practices.
Work-Related Limitations 3 The student mentions employment as a limiting factor in course participation.
Invested Effort 3 The student highlights having invested significant effort into the course.
Preference for Independent Work 2 The student prefers to work alone rather than in groups.
Deadline-Driven Approach 2 The student operates primarily based on deadlines.
Lecturer-Induced Feelings of Inferiority 2 The student recounts an instance where the lecturer caused feelings of inadequacy.
Random Study Practices 2 The student indicates randomness in their study methods.
Course Workload for Working Students 2 The student highlights the demanding nature of the course workload for employed students.
Dependence on Instructors 1 The student’s progress depends heavily on instructor guidance and assistance.
Burnout Due to Workload 1 The student expresses burnout resulting from the course workload.
AI Tools 1 The student mentions using AI tools, such as ChatGPT, for assistance.
Lack of Motivation 1 The student expresses a lack of motivation regarding the course.
Language Barrier Challenges 1 The student identifies the language barrier as a challenge during the course.
Library Documentation 1 The student reports using documentation from the Jypeli library for project work.
Establishing Routine 1 The student demonstrates self-regulation by creating a routine for course completion.
Increasing Independence Over Time 1 The student notes enhanced self-regulation as the course progresses.
Expressed Motivation 1 The student conveys motivation toward the course.
Deliberate Note-Taking 1 The student practices self-regulation through conscious note-taking to enhance learning.
Low Goal Setting 1 The student sets low personal goals for the course.
Social Aspects of the Course 1 The student brings up matters related to social interactions in the course.
Challenge of Transitioning to University 1 The student mentions difficulties in transitioning to university-level studies.
Minimal Effort Invested 1 The student notes limited personal investment in the course.
Concern Over Effort Sufficiency 1 The student expresses concern about the adequacy of their effort in the course.
Adjusting Habits to Course Schedule 1 The student adapts their habits to align with the course timetable.
Awareness of Personal Goals 1 The student acknowledges and articulates personal goals for the course.
Table 3. Themes identified from content analysis of 29 student questionnaires to allow easier comparison and summary of student profiles. For each theme, theme name, number of student in which the theme emerged, and a brief description of the theme is provided.